# An Introduction to Single-Antenna Radio Astronomical Polarimetry

### v2

@@ 1 Introduction
Before embarking on this approach, Section 2 of this paper reviews the geometric description of polarization introduced by Stokes (1852), beginning with an ideal, monochromatic source of electromagnetic radiation. In Section 3, the Stokes parameters are related to the second-order statistics of the electric field as represented by the coherency matrix. This connection is further explored in Section 4, where linear transformations of the electric field are related to Euclidean rotations and Lorentz boosts of the Stokes parameters. In Section 5, these transformations are applied to a decomposition of the signal path that extends the Hamaker et al. (1996) model to include the effects of reflection and down-conversion. Inverting such transformations to recover the original polarization state is discussed in Section 6, which includes a first-order calibration solution based on the ideal feed assumptions. Section 7 introduces some criteria for selecting a more complete model of the instrumental response, and Section 8 concludes with some pointers to further reading.
@@ 2 Monochromatic Light
Figure 1: A left-handed ellipse is traced by the electric field vector on each cycle of a monochromatic wave travelling in the positive $z$ direction, as observed when looking toward the source. The position angle, $\pi/2\leq\psi\leq\pi/2$, describes the orientation of the semi-major axis of the ellipse with respect to the $x$-axis; it is positive when the $x^{\prime}$ and $y^{\prime}$ axes are rotated in a counter-clockwise direction with respect to the $x$ and $y$ axes. The tangent of the ellipticity angle, $\pi/4\leq\chi\leq\pi/4$, is equal to the ratio between the semi-minor axis and the semi-major axis; it is positive for a left-hand polarized wave and negative for a right-hand polarized wave. The size of the ellipse is defined by $r$, which is equal to the length of the hypotenuse of the right triangle that contains $\chi$.
@@
When $\chi=0$, $\epsilon_{y}^{\prime}=0$ and the electric field vector oscillates along a line defined by the $x^{\prime}$ axis, which is oriented with respect to the $x$-axis by $\psi$. When $\chi>0$, the electric field travels in a clockwise direction about the ellipse, as seen by an observer looking toward the source, which is defined as a left-hand polarized wave (IEEE, 1983). When $\chi<0$, the counter-clockwise traverse of the electric field vector is defined as right-hand polarization. In the special case that $\chi=\pm\pi/4$, the amplitudes of $\epsilon_{x}^{\prime}$ and $\epsilon_{y}^{\prime}$ are equal and the electric field traces a circle in the xxyy plane. In general, polarization state depends on both the amplitudes of the electric field components and the relative phase between them.
@@ 3.1 Stokes Parameters
where $I$ is the intensity of the light, $p$ is the degree of polarization (see Section 3.2), and $\chi$ and $\psi$ are the ellipticity and orientation angles that define the polarization ellipse.
@@
is separable into scalar and vector components, $S_{0},\underline{\mbox{$S$}}]$, where $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})$ is the polarization vector with length $|\underline{\mbox{$S$}}|=Ip$ and direction defined by $2\psi$ and $2\chi$, as depicted in Figure 2.
Figure 2: The spherical coordinates of the Stokes polarization vector, $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})=(Q,U,V)$, include the vector length $|\underline{\mbox{$S$}}|=Ip$, longitude $2\psi$, and latitude $2\chi$. Points in the $S_{1}$$S_{2}$ plane ($S_{3}=0$) represent linearly polarized states, points above this plane ($S_{3}>0$) represent left-hand elliptically polarized states and the positive $S_{3}$ pole represents left-hand circular polarization.
@@ 3.4 Polarization Ellipse
Stokes $\pm Q$: If $a_{y}=0$ in equation (1), then the radiation is 100% linearly polarized with the electric field vector oscillating along the x-axis. In this case, $\chi=0$, $\psi=0$; Stokes $Q=S_{1}=S_{0}$ is positive; and Stokes $S_{2}=S_{3}=0$.
If $a_{x}=0$, then the radiation is 100% linearly polarized with electric field vector oscillating along the y-axis. In this case, $\chi=0$, $\psi=\pm\pi/2$; Stokes $Q=S_{1}=-S_{0}$ is negative; and Stokes $S_{2}=S_{3}=0$.
Stokes $\pm U$: If $a_{x}=a_{y}$ and the relative phase difference $\Delta\phi=\phi_{x}-\phi_{y}=0$, then $\epsilon_{x}(t)=\epsilon_{y}(t)$ and the radiation is 100% linearly polarized with the electric field vector oscillating along the line defined by $y=x$ (i.e., offset by $45^{\circ}$ with respect to the x-axis). In this case, $\chi=0$, $\psi=\pi/4$; Stokes $U=S_{2}=S_{0}$ is positive; and Stokes $S_{1}=S_{3}=0$.
If $a_{x}=a_{y}$, and $\Delta\phi=\pm\pi$, then $\epsilon_{x}(t)=-\epsilon_{y}(t)$ and the radiation is 100% linearly polarized with the electric field vector oscillating along the line defined by y=xy=-x (i.e., offset by $45^{\circ}$ with respect to the x-axis). In this case, $\chi=0$, $\psi=-\pi/4$; Stokes $U=S_{2}=-S_{0}$ is negative; and Stokes $S_{1}=S_{3}=0$.
Stokes $\pm V$: If $a_{x}=a_{y}$ and $\Delta\phi=-\pi/2$, then the phase of $\epsilon_{y}$ leads that of $\epsilon_{x}$ by $90^{\circ}$ and the radiation is 100% circularly polarized with the electric field vector tracing a clockwise circle in the x-y plane, as seen by an observer looking toward the source; this is defined as a left-hand circularly polarized (LCP) wave. In this case, $\chi=\pi/4$, $\psi$ is undefined; Stokes $V=S_{3}=S_{0}$ is positive; and Stokes $S_{1}=S_{2}=0$.
If $a_{x}=a_{y}$ and $\Delta\phi=\pi/2$, then the phase of $\epsilon_{x}$ leads that of $\epsilon_{y}$ by $90^{\circ}$ and the radiation is 100% circularly polarized with the electric field vector tracing a counter-clockwise circle in the xy plane, as seen by an observer looking toward the source; this is defined as a right-hand circularly polarized (RCP) wave. In this case, $\chi=-\pi/4$, $\psi$ is undefined; Stokes $V=S_{3}=-S_{0}$ is negative; and Stokes $S_{1}=S_{2}=0$.
Figure 3: The polarization ellipses inscribed by the electric field vector for the six special cases of monochromatic light described in Section 3.4.
3.5 Adopted Standard
@@ 3.6 Coherency Matrix
$S_{\mu}={\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\mbox{$\rho$}}.$ (16)
Proof 3.1.
${\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\mbox{$\rho$}}$ $=\frac{1}{2}S_{\kappa}\,{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\kappa}}$ Eqn. (14) $=S_{\kappa}\delta_{\mu\kappa}$ Eqn. (C4) $=S_{\mu}$
@@
Example: Consider $\mu=3$, $S_{3}=\mbox{$\sigma$}_{3}\mbox{$:$}\mbox{$\rho$}$ Eqn. (16) $=\mbox{$\sigma$}_{3}\mbox{$:$}\langle\mbox{$e$}\otimes\mbox{$e$}^{\dagger}\rangle$ Eqn. (13) $={\rm Tr}\left[{\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\left(\begin{array}[]{cc}\langle e_{0}e_{0}^{*}\rangle&\langle e_{0}e_{1}^{*}\rangle\\ \langle e_{1}e_{0}^{*}\rangle&\langle e_{1}e_{1}^{*}\rangle\end{array}\right)}\right]$ Eqn. (B15) $=-i\langle e_{1}e_{0}^{*}\rangle+i\langle e_{0}e_{1}^{*}\rangle$ $=2\,{\rm Im}[\langle e_{0}^{*}\,e_{1}\rangle].$
3.7 Invariant Interval
@@
Any polarization state that fails to satisfy equation (23) may be called invalid, nonphysical, and/or over-polarized. Although nonphysical, over-polarized states can arise during numerical analysis. For example, they can be produced through transformation by an impure Mueller matrix (e.g. Section 7.4). They can also arise when all four Stokes parameters are allowed to vary independently during modeling.
@@
$\tilde{s}_{\mu}\equiv\mbox{$\sigma$}_{\mu}:\tilde{\mbox{$\rho$}}\text{ \quad where \quad}\tilde{\mbox{$\rho$}}\equiv\mbox{$e$}\otimes\mbox{$e$}^{\dagger},$ (24)
@@ 3.8 Orthogonal Polarizations
${\tilde{\mbox{$\rho$}}_{a}}\,\mbox{$:$}\,{\tilde{\mbox{$\rho$}}_{b}}=|\mbox{$e$}_{a}^{\dagger}\cdot\mbox{$e$}_{b}|^{2}.$ (30)
Proof $3.$
@@
$\tilde{s}_{a,\mu}=\mbox{$\sigma$}_{\mu}:\tilde{\mbox{$\rho$}}_{a}\quad\text{and}\quad\tilde{s}_{b,\mu}=\mbox{$\sigma$}_{\mu}:\tilde{\mbox{$\rho$}}_{b},$
@@
${\tilde{\mbox{$\rho$}}_{a}}\,\mbox{$:$}\,{\tilde{\mbox{$\rho$}}_{b}}=\frac{1}{2}\left(\tilde{s}_{a,0}\tilde{s}_{b,0}+\tilde{\underline{\mbox{$s$}}}_{a}\cdot\tilde{\underline{\mbox{$s$}}}_{b}\right).$ (31)
Proof 3.4.
${\tilde{\mbox{$\rho$}}_{a}}$ $:{\tilde{\mbox{$\rho$}}_{b}}=(\tilde{s}_{a,\mu}\,\mbox{$\sigma$}_{\mu}/2):(\tilde{s}_{b,\nu}\,\mbox{$\sigma$}_{\nu}/2)$ $=\frac{1}{4}\tilde{s}_{a,\mu}\,\tilde{s}_{b,\nu}\,\mbox{$\sigma$}_{\mu}:\mbox{$\sigma$}_{\nu}$ $=\frac{1}{2}\tilde{s}_{a,\mu}\,\tilde{s}_{b,\nu}\,\delta_{\mu\nu}$ Eqn. (C4) $=\frac{1}{2}\tilde{s}_{a,\mu}\,\tilde{s}_{b,\mu}$
@@
4 Linear Transformations
Polarimetric studies require modeling and correcting the manner in which the measurement apparatus alters the incident radiation. In radio astronomy, the instrument includes everything from the antenna (including any reflectors involved in focusing and redirecting the radio waves) to the signal processing system used to compute the Stokes parameters. It is also necessary to model and correct Faraday rotation, which occurs in the Earth’s ionosphere and the interstellar medium. In this section, the response of a system is described using linear transformations of the electric field and the equivalent linear transformations of the Stokes parameters.
4.1 Jones Matrices
In the narrow-band (or quasi-monochromatic) approximation of an electromagnetic wave (see Appendix A.2), the response of a single receptor is defined by a complex-valued row vector, $\mbox{$r$}=(r_{0}\;\;r_{1})$, such that the voltage induced in the receptor by the incident electric field $\mbox{$e$}(t)$ is given by the scalar product, $v(t)=\mbox{$r$}\cdot\mbox{$e$}(t)$. The intensity of the response, $I=\langle v^{2}(t)\rangle$, is maximum when the polarization of the incident wave matches that of the receptor, and reduces to zero when the incident wave is orthogonally polarized.
@@ 4.1 Jones Matrices
These transformations are depicted in Figure 4 and defined and discussed in more detail in Section 5. Note that Figure 4 includes an additional phase convention correction $\underline{\underline{{\bf\Phi}}}$ that cannot be represented using a Jones matrix and therefore does not appear in equation (36). The phase convention correction can be represented by a Mueller matrix and is included in equation (42) of Section 4.3.
@@ 4.2 Congruence Transformations
As discussed in more detail in Section 4.5, the congruence transformation is a special case of the more general radio interferometer measurement equation for a single point source (Hamaker, 2000; Smirnov, 2011a). For a single antenna, the coherency matrix is insensitive to the absolute phase of $\bf J$}.
Proof 4.2.
@@
(The matrices in this equation are defined and discussed in more detail in Section 5.)

### v4

@@ I Introduction
Before embarking on this approach, Section II of this paper reviews the geometric description of polarization introduced by Stokes (1852), beginning with an ideal, monochromatic source of electromagnetic radiation. In Section III, the Stokes parameters are related to the second-order statistics of the electric field as represented by the coherency matrix. This connection is further explored in Section IV, where linear transformations of the electric field are related to Euclidean rotations and Lorentz boosts of the Stokes parameters. In Section V, these transformations are applied to a decomposition of the signal path that extends the Hamaker et al. (1996) model to include the effects of reflection and down-conversion. Inverting such transformations to recover the original polarization state is discussed in Section VI, which includes a first-order calibration solution based on the ideal feed assumptions. Section VII introduces some criteria for selecting a more complete model of the instrumental response, and Section VIII concludes with some pointers to further reading.
@@ II Monochromatic Light
Figure 1: A left-handed ellipse is traced by the electric field vector on each cycle of a monochromatic wave travelling in the positive $z$ direction, as observed when looking toward the source. The position angle, $\pi/2\leq\psi\leq\pi/2$, describes the orientation of the semi-major axis of the ellipse with respect to the $x$ axis; it is positive when the $x^{\prime}$ and $y^{\prime}$ axes are rotated in a counter-clockwise direction with respect to the $x$ and $y$ axes. The tangent of the ellipticity angle, $\pi/4\leq\chi\leq\pi/4$, is equal to the ratio between the semi-minor axis and the semi-major axis; it is positive for a left-hand polarized wave and negative for a right-hand polarized wave. The size of the ellipse is defined by $r$, which is equal to the length of the hypotenuse of the right triangle that contains $\chi$.
@@
When $\chi=0$, $\epsilon_{y}^{\prime}=0$ and the electric field vector oscillates along a line defined by the $x^{\prime}$ axis, which is oriented with respect to the $x$ axis by $\psi$. When $\chi>0$, the electric field travels in a clockwise direction about the ellipse, as seen by an observer looking toward the source, which is defined as a left-hand polarized wave (IEEE, 1983). When $\chi<0$, the counter-clockwise traverse of the electric field vector is defined as right-hand polarization. In the special cases of $\chi=\pm\pi/4$, the amplitudes of $\epsilon_{x}^{\prime}$ and $\epsilon_{y}^{\prime}$ are equal and the electric field traces a circle in the xxyy plane. In general, polarization state depends on both the amplitudes of the electric field components and the relative phase between them.
@@ III.1 Stokes Parameters
where $I$ is the intensity of the light, $p$ is the degree of polarization (see Section III.2), and $\chi$ and $\psi$ are the ellipticity and orientation angles that define the polarization ellipse.
@@
is separable into scalar and vector components, $S_{0},\underline{\mbox{$S$}}]$, where $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})^{T}$ is the polarization vector with length $|\underline{\mbox{$S$}}|=Ip$ and direction defined by $2\psi$ and $2\chi$, as depicted in Figure 2.
Figure 2: The spherical coordinates of the polarization vector, $\underline{\mbox{$S$}}=(S_{1},S_{2},S_{3})^{T}=(Q,U,V)^{T}$, include the vector length $|\underline{\mbox{$S$}}|=Ip$, longitude $2\psi$, and latitude $2\chi$. Points in the $S_{1}$$S_{2}$ plane ($S_{3}=0$) represent linearly polarized states, points above this plane ($S_{3}>0$) represent left-hand elliptically polarized states and the positive $S_{3}$ pole represents left-hand circular polarization.
@@ III.4 Polarization Ellipse
Stokes $\pm Q$: If $a_{y}=0$ in Equation (1), then the radiation is 100% linearly polarized with the electric field vector oscillating along the $x$ axis. In this case, $\chi=0$, $\psi=0$, Stokes $Q=S_{1}=S_{0}$ is positive, and Stokes $S_{2}=S_{3}=0$.
If $a_{x}=0$, then the radiation is 100% linearly polarized with electric field vector oscillating along the $y$ axis. In this case, $\chi=0$, $\psi=\pm\pi/2$, Stokes $Q=S_{1}=-S_{0}$ is negative, and Stokes $S_{2}=S_{3}=0$.
Stokes $\pm U$: If $a_{x}=a_{y}$ and the relative phase difference, $\Delta\phi=\phi_{x}-\phi_{y}=0$, then $\epsilon_{x}(t)=\epsilon_{y}(t)$ and the radiation is 100% linearly polarized with the electric field vector oscillating along the line defined by $y=x$ (i.e., offset by $45^{\circ}$ with respect to the $x$ axis). In this case, $\chi=0$, $\psi=\pi/4$, Stokes $U=S_{2}=S_{0}$ is positive, and Stokes $S_{1}=S_{3}=0$.
If $a_{x}=a_{y}$, and $\Delta\phi=\pm\pi$, then $\epsilon_{x}(t)=-\epsilon_{y}(t)$ and the radiation is 100% linearly polarized with the electric field vector oscillating along the line defined by y=xy=-x (i.e., offset by $45^{\circ}$ with respect to the $x$ axis). In this case, $\chi=0$, $\psi=-\pi/4$, Stokes $U=S_{2}=-S_{0}$ is negative, and Stokes $S_{1}=S_{3}=0$.
Stokes $\pm V$: If $a_{x}=a_{y}$ and $\Delta\phi=-\pi/2$, then the phase of $\epsilon_{y}$ leads that of $\epsilon_{x}$ by $90^{\circ}$ and the radiation is 100% circularly polarized with the electric field vector tracing a clockwise circle in the x-y plane, as seen by an observer looking toward the source; this is defined as a left-hand circularly polarized (LCP) wave. In this case, $\chi=\pi/4$, $\psi$ is undefined, Stokes $V=S_{3}=S_{0}$ is positive, and Stokes $S_{1}=S_{2}=0$.
If $a_{x}=a_{y}$ and $\Delta\phi=\pi/2$, then the phase of $\epsilon_{x}$ leads that of $\epsilon_{y}$ by $90^{\circ}$ and the radiation is 100% circularly polarized with the electric field vector tracing a counter-clockwise circle in the x-y plane, as seen by an observer looking toward the source; this is defined as a right-hand circularly polarized (RCP) wave. In this case, $\chi=-\pi/4$, $\psi$ is undefined, Stokes $V=S_{3}=-S_{0}$ is negative, and Stokes $S_{1}=S_{2}=0$.
Figure 3: The polarization ellipses inscribed by the electric field vector for the six special cases of monochromatic light described in Section III.4.
III.5 Adopted Standard
@@ III.6 Coherency Matrix
$S_{\mu}={\mbox{$\sigma$}_{\mu}}\mathbin{:}{\mbox{$\rho$}}.$ (16)
Proof III.1.
${\mbox{$\sigma$}_{\mu}}\mathbin{:}{\mbox{$\rho$}}$ $=\frac{1}{2}S_{\kappa}\,{\mbox{$\sigma$}_{\mu}}\mathbin{:}{\mbox{$\sigma$}_{\kappa}}$ Eqn. (14) $=S_{\kappa}\delta_{\mu\kappa}$ Eqn. (C4) $=S_{\mu}$
@@
Example: Consider $\mu=3$, $S_{3}=\mbox{$\sigma$}_{3}\mathbin{:}\mbox{$\rho$}$ Eqn. (16) $=\mbox{$\sigma$}_{3}\mathbin{:}\langle\mbox{$e$}\otimes\mbox{$e$}^{\dagger}\rangle$ Eqn. (13) $={\rm Tr}\left[{\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\left(\begin{array}[]{cc}\langle e_{0}e_{0}^{*}\rangle&\langle e_{0}e_{1}^{*}\rangle\\ \langle e_{1}e_{0}^{*}\rangle&\langle e_{1}e_{1}^{*}\rangle\end{array}\right)}\right]$ Eqn. (B15) $=-i\langle e_{1}e_{0}^{*}\rangle+i\langle e_{0}e_{1}^{*}\rangle$ $=2\,{\rm Im}[\langle e_{0}^{*}\,e_{1}\rangle].$
III.7 Invariant Interval
@@
Any polarization state that fails to satisfy Equation (23) may be called invalid, nonphysical, and/or over-polarized. Although nonphysical, over-polarized states can arise during numerical analysis. For example, they can be produced through transformation by an impure Mueller matrix (e.g. Section VII.4). They can also arise when all four Stokes parameters are allowed to vary independently during modeling.
@@
$\tilde{s}_{\mu}\equiv\mbox{$\sigma$}_{\mu}\mathbin{:}\tilde{\mbox{$\rho$}}\text{ \quad where \quad}\tilde{\mbox{$\rho$}}\equiv\mbox{$e$}\otimes\mbox{$e$}^{\dagger},$ (24)
@@ III.8 Orthogonal Polarizations
${\tilde{\mbox{$\rho$}}_{a}}\mathbin{:}{\tilde{\mbox{$\rho$}}_{b}}=|\mbox{$e$}_{a}^{\dagger}\cdot\mbox{$e$}_{b}|^{2}.$ (30)
Proof III.3.
@@
$\tilde{s}_{a,\mu}=\mbox{$\sigma$}_{\mu}\mathbin{:}\tilde{\mbox{$\rho$}}_{a}\quad\text{and}\quad\tilde{s}_{b,\mu}=\mbox{$\sigma$}_{\mu}\mathbin{:}\tilde{\mbox{$\rho$}}_{b},$
@@
${\tilde{\mbox{$\rho$}}_{a}}\mathbin{:}{\tilde{\mbox{$\rho$}}_{b}}=\frac{1}{2}\left(\tilde{s}_{a,0}\tilde{s}_{b,0}+\tilde{\underline{\mbox{$s$}}}_{a}\cdot\tilde{\underline{\mbox{$s$}}}_{b}\right).$ (31)
Proof III.4.
${\tilde{\mbox{$\rho$}}_{a}}$ $\mathbin{:}{\tilde{\mbox{$\rho$}}_{b}}=(\tilde{s}_{a,\mu}\,\mbox{$\sigma$}_{\mu}/2)\mathbin{:}(\tilde{s}_{b,\nu}\,\mbox{$\sigma$}_{\nu}/2)$ $=\frac{1}{4}\tilde{s}_{a,\mu}\,\tilde{s}_{b,\nu}\,\mbox{$\sigma$}_{\mu}\mathbin{:}\mbox{$\sigma$}_{\nu}$ $=\frac{1}{2}\tilde{s}_{a,\mu}\,\tilde{s}_{b,\nu}\,\delta_{\mu\nu}$ Eqn. (C4) $=\frac{1}{2}\tilde{s}_{a,\mu}\,\tilde{s}_{b,\mu}$
@@
IV Linear Transformations
Polarimetric studies require modeling and correcting the manner in which the measurement apparatus alter the incident radiation. In radio astronomy, the instrument includes everything from the antenna (including any reflectors involved in focusing and redirecting the radio waves) to the signal processing system used to compute the Stokes parameters. It is also necessary to model and correct Faraday rotation, which occurs in the Earth’s ionosphere and the interstellar medium. In this section, the response of a system is described using linear transformations of the electric field and the equivalent linear transformations of the Stokes parameters.
IV.1 Jones Matrices
In the narrow-band (or quasi-monochromatic) approximation of an electromagnetic wave (see Appendix A.2), the response of a single receptor is defined by a complex-valued row vector, $\mbox{$r$}=(r_{0},r_{1})$, such that the voltage induced in the receptor by the incident electric field $\mbox{$e$}(t)$ is given by the scalar product, $v(t)=\mbox{$r$}\cdot\mbox{$e$}(t)$. The intensity of the response, $I=\langle v^{2}(t)\rangle$, is maximum when the polarization of the incident wave matches that of the receptor, and reduces to zero when the incident wave is orthogonally polarized.
@@ IV.1 Jones Matrices
These transformations are depicted in Figure 4 and defined and discussed in more detail in Section V. Note that Figure 4 includes an additional phase convention correction $\underline{\underline{{\bf\Phi}}}$ that cannot be represented using a Jones matrix and therefore does not appear in Equation (36). The phase convention correction can be represented by a Mueller matrix and is included in Equation (42) of Section IV.3.
@@ IV.2 Congruence Transformations
As discussed in more detail in Section IV.5, the congruence transformation is a special case of the more general radio interferometer measurement equation for a single point source (Hamaker, 2000; Smirnov, 2011a). For a single antenna, the coherency matrix is insensitive to the absolute phase of $\bf J$}.
Proof IV.2.
@@
(The matrices in this equation are defined and discussed in more detail in Section V.)